# Moser’s twist theorem revisited

### v2

@@ 3.2. Criteria
$\Lambda_{\iota}(\kappa):=\max\left\{\sup_{v\in(\kappa)_{\iota}}\frac{\lambda(v)}{\Theta(v)},\sup_{v\in(\kappa)_{\iota}}\frac{\Theta(v)}{\lambda(v)}\right\},\quad\iota\in\{I,II\}.$
@@
Obviously, we only need to consider the case with $\kappa$ large enough. For $\kappa$ large enough, if $v=v^{\prime}+v^{\prime\prime}$, then we have $\lambda(v)=\lambda(v^{\prime})+\lambda(v^{\prime\prime})$ and $\Theta(v)=\Theta(v^{\prime})+\Theta(v^{\prime})$. It follows that $\frac{\lambda(v)}{\Theta(v)}$ (resp. $\frac{\Theta(v)}{\lambda(v)}$) is a convex combination of $\frac{\lambda(v^{\prime})}{\Theta(v^{\prime})}$ and $\frac{\lambda(v^{\prime\prime})}{\Theta(v^{\prime\prime})}$ (resp. $\frac{\Theta(v^{\prime})}{\lambda(v^{\prime})}$ and $\frac{\Theta(v^{\prime\prime})}{\lambda(v^{\prime\prime})}$). Hence, $\Lambda_{I}(\kappa)$ is nondecreasing (up to a subsequence). It suffices to verify the boundedness of $\{\Lambda_{I}(\kappa)\}_{\kappa\in S}$. Moreover, one can show that $\{\Lambda_{I}(\kappa)\}_{\kappa\in S}$ is even controlled by $\{\Lambda_{II}(\kappa)\}_{\kappa\in\mathbb{N}}$, for which one needs to show the following
@@
For each $v\in(\kappa)_{I}$ with $\kappa\in S$, there exist $v^{\prime}\in(\kappa^{\prime})_{II}$, $v^{\prime\prime}\in(\kappa^{\prime\prime})_{II}$ and a constant $C$ independent of $\kappa,\kappa^{\prime},\kappa^{\prime\prime}$ and such that
$\frac{1}{C}\frac{\lambda(v^{\prime\prime})}{\Theta(v^{\prime\prime})}\leq\frac{\lambda(v)}{\Theta(v)}\leq C\frac{\lambda(v^{\prime})}{\Theta(v^{\prime})}.$
Proof   Fix $n\in\mathbb{N}$. There holds (see for instance [7, V. (7.6)])
@@
$2^{j}\|q_{n+2^{j}}\alpha\|<\|q_{n}\alpha\|<(A_{\alpha}+1)^{2j}\|q_{n+2^{j}}\alpha\|.$
@@
$\|q_{n+2^{4}}\alpha\|<\frac{1}{2^{4}}\|q_{n}\alpha\|<\frac{1}{2^{\kappa+4}}<\frac{1}{2^{\kappa}}<2^{4}\|q_{n}\alpha\|<\|q_{n-2^{4}}\alpha\|<(A_{\alpha}+1)^{8}\|q_{n}\alpha\|.$
@@
  • Proof of $\frac{\lambda(v)}{\Theta(v)}\leq C\frac{\lambda(v^{\prime})}{\Theta(v^{\prime})}$.
Given $v\in(\kappa)_{I}$ with $\kappa=\phi(n)$, if $\lambda(v)\in[\|q_{n}\alpha\|,\frac{1}{2^{\kappa}}]$, then we only need to choose $v^{\prime}\in(\kappa)_{II}$ satisfying $\lambda(v^{\prime})=\|q_{n}\alpha\|$ such that
$\frac{\lambda(v)}{\Theta(v)}\leq 2^{4}\frac{\|q_{n}\alpha\|}{\Theta(v^{\prime})}=2^{4}\frac{\lambda(v^{\prime})}{\Theta(v^{\prime})}.$
In fact, by the orientation-preserving property (see Proposition 2.1), $\Theta(v^{\prime})<\Theta(v)$.
Otherwise, if $\lambda(v)\in[\frac{1}{2^{\kappa+4}},\|q_{n}\alpha\|)$, one choose $v^{\prime}\in(\kappa^{\prime})_{II}$ with $\kappa^{\prime}=\phi(n+2^{4})$ satisfying $\lambda(v^{\prime})=\|q_{n+2^{4}}\alpha\|$. Then we still have $\Theta(v^{\prime})<\Theta(v)$ from the orientation-preserving property. Moreover,
$\frac{\lambda(v)}{\Theta(v)}\leq\frac{\|q_{n}\alpha\|}{\Theta(v^{\prime})}<(A_{\alpha}+1)^{8}\frac{\|q_{n+2^{4}}\alpha\|}{\Theta(v^{\prime})}\leq(A_{\alpha}+1)^{8}\frac{\lambda(v^{\prime})}{\Theta(v^{\prime})}.$
  • Proof of $\frac{1}{C}\frac{\lambda(v^{\prime\prime})}{\Theta(v^{\prime\prime})}\leq\frac{\lambda(v)}{\Theta(v)}$.
Given $v\in(\kappa)_{I}$ with $\kappa=\phi(n)$, we only need to choose $v^{\prime\prime}\in(\kappa^{\prime\prime})_{II}$ with $\kappa^{\prime\prime}=\phi(n-2^{4})$ satisfying $\lambda(v^{\prime\prime})=\|q_{n-2^{4}}\alpha\|$. Similarly, by the orientation-preserving property (see Proposition 2.1), $\Theta(v^{\prime\prime})>\Theta(v)$. It follows that
$\frac{1}{2^{4}(A_{\alpha}+1)^{8}}\frac{\lambda(v^{\prime\prime})}{\Theta(v^{\prime\prime})}<\frac{1}{2^{4}}\frac{\|q_{n}\alpha\|}{\Theta(v^{\prime\prime})}\leq\frac{\lambda(v)}{\Theta(v)}.$
[…]
@@
If $v\in(\kappa_{0})_{II}$ with $\kappa_{0}=\phi(n_{0})\in S$, then $\Theta(v)\sim\lambda(v)$.
Proof   Note that $\lambda(v)\sim\|q_{n_{0}}\alpha\|\sim q_{n_{0}}^{-1}$. Based on (9) and (11), we write $\bar{N}:=\bar{N}(\kappa_{0})$. It follows from Proposition 3.1 that one can find $k_{0},k_{1}\in\{0,\ldots,q_{\bar{N}-1}\}$ such that
$\Theta(F^{k_{0}}(v))\gtrsim\lambda(v),\quad\Theta(F^{k_{1}}(v))\lesssim\lambda(v).$
@@
$\Theta(F^{k_{0}}(v))=\|x_{j+k_{0}}-x_{i+k_{0}}\|,\quad\Theta(F^{k_{1}}(v))=\|x_{j+k_{1}}-x_{i+k_{1}}\|.$
@@
$\left|\sum_{l=0}^{k_{1}-k_{0}-1}(y_{j+l}-y_{i+l})\right|\lesssim\frac{1}{q_{n_{0}}^{1+\frac{\varepsilon}{2}}},$
@@
$\Theta(v)=\|x_{j}-x_{i}\|\sim\frac{1}{q_{n_{0}}}\sim\lambda(v).$
[…]
@@ 3.3. Reduction
Let us recall $v$ connects $(x_{i},y_{i})$ and $(x_{j},y_{j})$. Let $s(v)$ be the first-order difference quotient along the minimal configuration, i.e. $s(v):=\frac{r(v)}{\Theta(v)}$ where $r(v):=y_{j}-y_{i}$. Based on the Aubry-Mather theory, we know that there exists a constant $s_{\max}$ such that $|s(v)|\leq s_{\max}$ ([6, Theorem 14.1]). Note that one can take $s_{\max}=2$ for $G_{n}$. Moreover, if
@@
(12) $\frac{1}{3}\leq\frac{1}{1+s_{\max}}\leq\frac{\Theta(F(v))}{\Theta(v)}\leq 1+s_{\max}\leq 3.$
We also note that $\frac{\Theta(F^{j}(v))}{\Theta(v)}$ behaves like a multiplicative cocycle. Inspired by the theory of the circle diffeomorphism, we consider the expansion
$K^{0}(j|v):=\ln\frac{\Theta(F^{j}(v))}{\Theta(v)}.$
Denote
@@
$\mathbb{K}_{N}^{0}(\kappa):=\sup\{|K^{0}(q_{N}|v)|:\ v\in(\kappa)_{II}\};$
[…]
@@
(13) $\tilde{\mathbb{K}}_{N}^{0}(\kappa)=\sup\{|K^{0}(j|v)|:\ -q_{N}\leq j\leq q_{N},\ \ v\in(\kappa)_{II}\}.$
@@
(14) $\tilde{\mathbb{K}}_{\bar{N}(\kappa)}^{0}(\kappa)\leq C,$
@@
Proof   We use the same strategy as the proof of Lemma 3.6. Note that $\lambda(v)\sim\|q_{n_{\kappa}}\alpha\|\sim q_{n_{\kappa}}^{-1}$. By Proposition 3.1, one can find $k_{0},k_{1}\in\{0,\ldots,q_{\bar{N}(\kappa)-1}\}$ such that
$\Theta(F^{k_{0}}(v))\gtrsim\lambda(v),\quad\Theta(F^{k_{1}}(v))\lesssim\lambda(v).$
@@
(15) $\Theta(F^{j}(v))\gtrsim\lambda(v)\ \ \mathrm{or}\ \ \Theta(F^{j}(v))\lesssim\lambda(v).$
@@
$e^{-C}\leq\frac{\Theta(F^{k_{1}}(v))}{\Theta(F^{j}(v))}\leq e^{C}.$
@@
$\Theta(F^{j}(v))\sim\Theta(F^{k_{1}}(v))\sim\lambda(v).$
[…]
@@
  • Criterion 3: The sequence $\{\tilde{\mathbb{K}}_{\bar{N}(\kappa)}^{0}(\kappa)\}_{\kappa\in\mathbb{N}}$ is bounded.
@@ 4.1. Distortions
Let $v_{1},v_{2}$ be two chords. Denote by $\lambda(v_{1},v_{2})$ the average of the $\lambda$-length of the span of the two vectors and that of the part between them. To verify Criterion 3, we only need to consider the case with $\lambda(v_{1},v_{2})$ small enough. We assume, once and for all, $v_{1}$ connects $(x_{i_{1}},y_{i_{1}})$ and $(x_{j_{1}},y_{j_{1}})$ and $v_{2}$ connects $(x_{i_{2}},y_{i_{2}})$ and $(x_{j_{2}},y_{j_{2}})$. Moreover, we require
@@
$\lambda(v_{1},v_{2})=\frac{1}{2}(\|(i_{2}-i_{1})\alpha\|+\|(j_{2}-j_{1})\alpha\|).$
@@
$\theta(v_{1},v_{2}):=x_{i_{2}}-x_{i_{1}},\quad\Theta(v_{1},v_{2}):=\|\theta(v_{1},v_{2})\|.$
@@
$\frac{1}{(A_{\alpha}+1)^{6}}q_{r-2\gamma_{0}}\leq|i_{2}-i_{1}|\leq(A_{\alpha}+1)^{6}q_{r-2\gamma_{0}},$
$\lambda(v_{1},v_{2})\in[\|q_{r-2\gamma_{0}}\alpha\|,2^{4}\|q_{r-2\gamma_{0}}\alpha\|],$
where $\gamma_{0}$ is defined by (9). It follows from (8) that
$\frac{2}{q_{r}}<\lambda(v_{1},v_{2})<\frac{1}{q_{r-2\gamma_{0}}}.$
@@
$\tilde{\mathbb{K}}_{N}^{1}(\kappa,r):=\sup\{|K^{1}(j|v_{1},v_{2})|:\ -q_{N}\leq j\leq q_{N},\ \ (v_{1},v_{2})\in(\kappa,r)_{II}\}.$
@@
For each $\kappa\in\mathbb{N}$, and each $N\leq n_{\kappa}:=\psi(\kappa)$, there holds
$\mathbb{K}_{N}^{0}(\kappa)\leq 2\tilde{\mathbb{K}}_{N}^{1}(\kappa,N).$

### v3

@@ 3.2. Criteria
$\Lambda_{\iota}(\kappa):=\max\left\{\sup_{v\in(\kappa)_{\iota}}\frac{\lambda_{1}(v)}{\Theta_{1}(v)},\sup_{v\in(\kappa)_{\iota}}\frac{\Theta_{1}(v)}{\lambda_{1}(v)}\right\},\quad\iota\in\{I,II\}.$
@@
Obviously, we only need to consider the case with $\kappa$ large enough. For $\kappa$ large enough, if $v=v^{\prime}+v^{\prime\prime}$, then we have $\lambda_{1}(v)=\lambda_{1}(v^{\prime})+\lambda_{1}(v^{\prime\prime})$ and $\Theta_{1}(v)=\Theta_{1}(v^{\prime})+\Theta_{1}(v^{\prime\prime})$. It follows that $\frac{\lambda_{1}(v)}{\Theta_{1}(v)}$ (resp. $\frac{\Theta_{1}(v)}{\lambda_{1}(v)}$) is a convex combination of $\frac{\lambda_{1}(v^{\prime})}{\Theta_{1}(v^{\prime})}$ and $\frac{\lambda_{1}(v^{\prime\prime})}{\Theta_{1}(v^{\prime\prime})}$ (resp. $\frac{\Theta_{1}(v^{\prime})}{\lambda_{1}(v^{\prime})}$ and $\frac{\Theta_{1}(v^{\prime\prime})}{\lambda_{1}(v^{\prime\prime})}$). Hence, $\Lambda_{I}(\kappa)$ is nondecreasing (up to a subsequence). It suffices to verify the boundedness of $\{\Lambda_{I}(\kappa)\}_{\kappa\in S}$. Moreover, one can show that $\{\Lambda_{I}(\kappa)\}_{\kappa\in S}$ is even controlled by $\{\Lambda_{II}(\kappa)\}_{\kappa\in\mathbb{N}}$, for which one needs to show the following
@@
For each $v\in(\kappa)_{I}$ with $\kappa\in S$, there exist $v^{\prime}\in(\kappa^{\prime})_{II}$ and $v^{\prime\prime}\in(\kappa^{\prime\prime})_{II}$ such that
$\frac{1}{2^{4}(A_{\alpha}+1)^{8}}\frac{\lambda_{1}(v^{\prime\prime})}{\Theta_{1}(v^{\prime\prime})}<\frac{\lambda_{1}(v)}{\Theta_{1}(v)}<(A_{\alpha}+1)^{8}\frac{\lambda_{1}(v^{\prime})}{\Theta_{1}(v^{\prime})},$
where $A_{\alpha}$ denotes an upper bound for all the partial quotients $a_{i}$ in the continued fraction expansion of $\alpha$.
Proof.
Fix $n\in\mathbb{N}$. There holds (see for instance [7, V. (7.6)])
@@
$2^{j}\|q_{n+2j}\alpha\|<\|q_{n}\alpha\|<(A_{\alpha}+1)^{2j}\|q_{n+2j}\alpha\|.$
@@
$\|q_{n+8}\alpha\|<\frac{1}{2^{4}}\|q_{n}\alpha\|<\frac{1}{2^{\kappa+4}}<\frac{1}{2^{\kappa}}<2^{4}\|q_{n}\alpha\|<\|q_{n-8}\alpha\|<(A_{\alpha}+1)^{8}\|q_{n}\alpha\|.$
@@
  • Proof of $\frac{\lambda_{1}(v)}{\Theta_{1}(v)}<(A_{\alpha}+1)^{8}\frac{\lambda_{1}(v^{\prime})}{\Theta_{1}(v^{\prime})}$.
Given $v\in(\kappa)_{I}$ with $\kappa=\phi(n)$, if $\lambda_{1}(v)\in[\|q_{n}\alpha\|,\frac{1}{2^{\kappa}}]$, then we only need to choose $v^{\prime}\in(\kappa)_{II}$ satisfying $\lambda_{1}(v^{\prime})=\|q_{n}\alpha\|$ such that
$\frac{\lambda_{1}(v)}{\Theta_{1}(v)}<2^{4}\frac{\|q_{n}\alpha\|}{\Theta_{1}(v^{\prime})}=2^{4}\frac{\lambda_{1}(v^{\prime})}{\Theta_{1}(v^{\prime})}.$
In fact, by the orientation-preserving property (see Proposition 2.1), $\Theta_{1}(v^{\prime})<\Theta_{1}(v)$.
Otherwise, if $\lambda_{1}(v)\in[\frac{1}{2^{\kappa+4}},\|q_{n}\alpha\|)$, one choose $v^{\prime}\in(\kappa^{\prime})_{II}$ with $\kappa^{\prime}=\phi(n+8)$ satisfying $\lambda_{1}(v^{\prime})=\|q_{n+8}\alpha\|$. Then we still have $\Theta_{1}(v^{\prime})<\Theta_{1}(v)$ from the orientation-preserving property. Moreover,
$\frac{\lambda_{1}(v)}{\Theta_{1}(v)}\leq\frac{\|q_{n}\alpha\|}{\Theta_{1}(v^{\prime})}<(A_{\alpha}+1)^{8}\frac{\|q_{n+8}\alpha\|}{\Theta_{1}(v^{\prime})}\leq(A_{\alpha}+1)^{8}\frac{\lambda_{1}(v^{\prime})}{\Theta_{1}(v^{\prime})}.$
  • Proof of $\frac{1}{2^{4}(A_{\alpha}+1)^{8}}\frac{\lambda_{1}(v^{\prime\prime})}{\Theta_{1}(v^{\prime\prime})}<\frac{\lambda_{1}(v)}{\Theta_{1}(v)}$.
Given $v\in(\kappa)_{I}$ with $\kappa=\phi(n)$, we only need to choose $v^{\prime\prime}\in(\kappa^{\prime\prime})_{II}$ with $\kappa^{\prime\prime}=\phi(n-8)$ satisfying $\lambda_{1}(v^{\prime\prime})=\|q_{n-8}\alpha\|$. Similarly, by the orientation-preserving property (see Proposition 2.1), $\Theta_{1}(v^{\prime\prime})>\Theta_{1}(v)$. It follows that
$\frac{1}{2^{4}(A_{\alpha}+1)^{8}}\frac{\lambda_{1}(v^{\prime\prime})}{\Theta_{1}(v^{\prime\prime})}<\frac{1}{2^{4}}\frac{\|q_{n}\alpha\|}{\Theta_{1}(v^{\prime\prime})}\leq\frac{\lambda_{1}(v)}{\Theta_{1}(v)}.$
This completes the proof of Lemma 3.6. ∎
@@
There exists a constant $N_{0}>0$ such that for any $n_{0}>N_{0}$, if $v\in(\kappa_{0})_{II}$ with $\kappa_{0}=\phi(n_{0})\in S$, then $\Theta_{1}(v)\sim\lambda_{1}(v)$. Namely, there exists a constant $C>0$, independent of $n$, such that
$\frac{1}{C}<\frac{\lambda_{1}(v)}{\Theta_{1}(v)}\leq C.$
Proof.
Note that $\lambda_{1}(v)\sim\|q_{n_{0}}\alpha\|\sim q_{n_{0}}^{-1}$. Based on (12) and (14), we write $\bar{N}:=\bar{N}(\kappa_{0})$. It follows from Proposition 3.2 that one can find $k_{0},k_{1}\in\{0,\ldots,q_{\bar{N}-1}\}$ such that
$\Theta_{1}(F^{k_{0}}(v))\gtrsim\lambda_{1}(v),\quad\Theta_{1}(F^{k_{1}}(v))\lesssim\lambda_{1}(v).$
@@
$\Theta_{1}(F^{k_{0}}(v))=\|x_{j+k_{0}}-x_{i+k_{0}}\|,\quad\Theta_{1}(F^{k_{1}}(v))=\|x_{j+k_{1}}-x_{i+k_{1}}\|.$
@@
(15) $\left|\sum_{l=0}^{k_{1}-k_{0}-1}(y_{j+l}-y_{i+l})\right|\lesssim\frac{1}{q_{n_{0}}^{1+\frac{\varepsilon}{2}}},$
@@
$\Theta_{1}(v)=\|x_{j}-x_{i}\|\sim\frac{1}{q_{n_{0}}}\sim\lambda_{1}(v).$
@@ 3.3. Reduction
Let us recall $v$ connects $(x_{i},y_{i})$ and $(x_{j},y_{j})$. Let $s(v)$ be the first-order difference quotient along the minimal configuration, i.e. $s(v):=\frac{r(v)}{\Theta_{1}(v)}$ where $r(v):=y_{j}-y_{i}$. Based on the Aubry-Mather theory, we know that there exists a constant $s_{\max}$ such that $|s(v)|\leq s_{\max}$ ([6, Theorem 14.1]). Note that one can take $s_{\max}=2$ for $G_{n}$. Moreover, if
@@
(16) $\frac{1}{3}\leq\frac{1}{1+s_{\max}}\leq\frac{\Theta_{1}(F(v))}{\Theta_{1}(v)}\leq 1+s_{\max}\leq 3.$
We also note that $\frac{\Theta_{1}(F^{j}(v))}{\Theta_{1}(v)}$ behaves like a multiplicative cocycle. Inspired by the theory of the circle diffeomorphism, we consider the expansion
$K^{0}(j|v):=\ln\frac{\Theta_{1}(F^{j}(v))}{\Theta_{1}(v)}.$
Denote for $\iota\in\{I,II\}$,
@@
$\mathbb{K}_{N}^{0}(\kappa):=\sup\{|K^{0}(q_{N}|v)|:\ v\in(\kappa)_{\iota}\};$
(17) $\widetilde{\mathbf{K}}_{N}^{0}(\kappa):=\sup\{|K^{0}(j|v)|:\ 0\leq j\leq q_{N},\ \ v\in(\kappa)_{\iota}\}.$
@@
(18) $\widetilde{\mathbf{K}}_{N}^{0}(\kappa)=\sup\{|K^{0}(j|v)|:\ -q_{N}\leq j\leq q_{N},\ \ v\in(\kappa)_{\iota}\}.$
@@
(19) $\widetilde{\mathbf{K}}_{\bar{N}(\kappa)}^{0}(\kappa)\leq C,$
@@
Proof.
We use the same strategy as the proof of Lemma 3.7. Note that for $\kappa$ sufficiently large such that $n_{\kappa}>N_{0}$, where $N_{0}$ is given by Lemma 3.7, we have $\lambda_{1}(v)\sim\|q_{n_{\kappa}}\alpha\|\sim q_{n_{\kappa}}^{-1}$. By Proposition 3.2, one can find $k_{0},k_{1}\in\{0,\ldots,q_{\bar{N}(\kappa)-1}\}$ such that
$\Theta_{1}(F^{k_{0}}(v))\gtrsim\lambda_{1}(v),\quad\Theta_{1}(F^{k_{1}}(v))\lesssim\lambda_{1}(v).$
@@
(20) $\Theta_{1}(F^{j}(v))\gtrsim\lambda_{1}(v)\ \ \mathrm{or}\ \ \Theta_{1}(F^{j}(v))\lesssim\lambda_{1}(v).$
@@
$e^{-C}\leq\frac{\Theta_{1}(F^{k_{1}}(v))}{\Theta_{1}(F^{j}(v))}\leq e^{C}.$
@@
$\Theta_{1}(F^{j}(v))\sim\Theta_{1}(F^{k_{1}}(v))\sim\lambda_{1}(v).$
For the second case in (20), we obtain the same result by comparing $\Theta_{1}(F^{j}(v))$ with $\Theta_{1}(F^{k_{0}}(v))$. In particular, we have $\Theta_{1}(v)\sim\lambda_{1}(v)$. This completes the proof. ∎
@@
  • Criterion 3: The sequence $\{\widetilde{\mathbf{K}}_{\bar{N}(\kappa)}^{0}(\kappa)\}_{\kappa\in\mathbb{N}}$ is bounded.
@@ 4.1. Distortions
Let $v_{1},v_{2}$ be two chords. Denote by $\lambda_{2}(v_{1},v_{2})$ the average of the $\lambda$-length of the span of the two chords and that of the part between them. To verify Criterion 3, we only need to consider the case with $\lambda_{2}(v_{1},v_{2})$ small enough. We assume, once and for all, $v_{1}$ connects $(x_{i_{1}},y_{i_{1}})$ and $(x_{j_{1}},y_{j_{1}})$ and $v_{2}$ connects $(x_{i_{2}},y_{i_{2}})$ and $(x_{j_{2}},y_{j_{2}})$. Moreover, we require
@@
$\lambda_{2}(v_{1},v_{2})=\frac{1}{2}(\|(i_{2}-i_{1})\alpha\|+\|(j_{2}-j_{1})\alpha\|).$
@@
$\theta(v_{1},v_{2}):=x_{i_{2}}-x_{i_{1}},\quad\Theta_{2}(v_{1},v_{2}):=\|\theta(v_{1},v_{2})\|.$
@@
$\frac{1}{q_{r+2}}\leq\lambda_{2}(v_{1},v_{2})\leq\frac{1}{q_{r}}.$
@@
$\widetilde{\mathbf{K}}_{N}^{1}(\kappa,r)=\sup\{|K^{1}(j|v_{1},v_{2})|:\ -q_{N}\leq j\leq q_{N},\ \ (v_{1},v_{2})\in(\kappa,r)_{II}\}.$
@@
For each $\kappa\in\mathbb{N}$, and each $N\leq\tilde{N}(\kappa)$, there holds
$\mathbb{K}_{N}^{0}(\kappa)\leq 2\widetilde{\mathbf{K}}_{N}^{1}(\kappa,N).$